# An analogue of Reed’s conjecture for digraphs

### v2

@@ 3 Colouring sparse digraphs
Our first goal is to generalise the following result due to Molloy and Reed [27, Theorem 10.5] on the chromatic number of sparse graphs. For a vertex $v$ of a graph $G$, we denote by $m(v)$ the number of edges of $G[N(v)]$. We say that $G$ is $B$-sparse if $m(v)\leq\binom{\Delta(G)}{2}-B$ for every $v\in V(G)$.
@@
$\chi(G)\leq(\Delta+1)-\frac{B}{e^{6}\Delta}.$
@@
There exists $\Delta_{1}\geq 2$ such that the following holds. Let $D$ be a digraph with maximum degree $\Delta_{\max}(D)=\Delta\geq\Delta_{1}$ and $a\in\mathbb{R}$ be such that $a>\frac{\log^{3}\Delta}{\Delta-1}$, and for every vertex $v\in V(D)$, $\min(m^{+}(v),m^{-}(v))\leq(1-a)\Delta(\Delta-1)$. Then
$\vec{\chi}(D)\leq\left(1-\frac{a}{5e^{7}}\right)\left(\Delta+1\right).$
@@
By assumption $D$ is $B_{a}$-sparse for $B_{a}=a\Delta(\Delta-1)>\Delta\log^{3}\Delta$. Hence, by taking $\Delta_{1}\geq\Delta_{0}$, we obtain by Theorem 1.7 that
$\vec{\chi}(D)\leq\left(1-\frac{a}{4e^{7}}\right)\left(\Delta+1\right)+\frac{a}{2e^{7}}\leq\left(1-\frac{a}{5e^{7}}\right)\left(\Delta+1\right),$
where in the second inequality we used $\Delta_{1}\geq 9$. ∎
@@
Let $D$ be a $B$-sparse digraph with maximum degree $\Delta_{\max}(D)=\Delta$. There exists a $B$-sparse $\Delta$-diregular digraph $H$ such that $\vec{\chi}(D)\leq\vec{\chi}(H)$.
@@
[…]
@@
|$\left|\{c\in[k]\mid\exists u,w\in N^{-}(v),u\neq w,\alpha(u)=\alpha(w)=c\}\right|\geq\ell$
or |$\left|\{c\in[k]\mid\exists u,w\in N^{+}(v),u\neq w,\alpha(u)=\alpha(w)=c\}\right|\geq\ell.$
@@
Let $D$ be a $\Delta$-diregular digraph and $k,\ell$ be two integers such that $k\leq\Delta+1-\ell$. If $D$ admits a partial $(k,\ell)$-dicolouring, then $\vec{\chi}(D)\leq\Delta+1-\ell$.
@@
Assume that $D$ admits a partial $(k,\ell)$-dicolouring. We let $\alpha$ be a partial $(\Delta+1-\ell,\ell)$-dicolouring of $D$ for which the number of uncoloured vertices is minimum, the existence of which is guaranteed as $k\leq\Delta+1-\ell$. We claim that $\alpha$ is a dicolouring of $D$, hence implying $\vec{\chi}(D)\leq\Delta+1-\ell$. Assume this is not the case, and there is a vertex $v$ which is not coloured by $\alpha$. By definition of a partial $(\Delta+1-\ell,\ell)$-dicolouring, we have $|\alpha(N^{-}(v))|\leq\Delta-\ell$ or $|\alpha(N^{+}(v))|\leq\Delta-\ell$. Hence we can choose for $v$ a colour of $[\Delta+1-\ell]$ that is not appearing either in its in- or in its out-neighbourhood. Colouring $v$ with this specific colour extends $\alpha$ into a partial $(\Delta+1-\ell,\ell)$-dicolouring of $D$ with fewer uncoloured vertices, a contradiction to the choice of $\alpha$. ∎
@@
We do not give the explicit value of $\Delta_{0}$, we simply assume in what follows that it is arbitrarily large. We assume that $D$ is $\Delta$-diregular, for otherwise we may prove the result for the $B$-sparse $\Delta$-diregular digraph $H$ obtained by Lemma 3.3, and conclude that $\vec{\chi}(D)\leq\vec{\chi}(H)\leq(\Delta+1)-\frac{B}{4e^{7}\Delta}$.
We use the probabilistic method. Let us fix $k=\left\lfloor\frac{\Delta}{2}\right\rfloor$. We randomly colour every vertex of $D$ with a colour of $[k]$. Then we simultaneously uncolour every vertex $v$ having at least one in-neighbour and at least one out-neighbour sharing its colour. Our goal is to show that with non-zero probability, the obtained colouring $\alpha$ is a partial $(k,\ell)$-dicolouring of $D$, where $\ell=\frac{B}{4e^{7}\Delta}$. Observe that we have $k\leq\Delta+1-\ell$ as $B\leq\Delta(\Delta-1)$. Hence, if $\alpha$ is indeed a partial $(k,\ell)$-dicolouring of $D$, the result follows from Lemma 3.4. Note that $\alpha$ is necessarily a dicolouring of an induced subdigraph of $D$, for otherwise $D$, partially coloured with $\alpha$, contains a monochromatic directed cycle, the vertices of which should have been uncoloured in the process. It remains to show that, for every vertex of $D$, $\ell$ colours are repeated either in its in-neighbourhood or in its out-neighbourhood.
For each vertex $v$, we let $N_{v}$ be $N^{+}(v)$ if $m^{+}(v)\leq m^{-}(v)$ and $N^{-}(v)$ otherwise. We also let $B_{v}$ be the number of missing arcs in $D[N_{v}]$, that is $B_{v}=\Delta(\Delta-1)-\min(m^{-}(v),m^{+}(v))$. Note that $B_{v}\geq B>\Delta\log^{3}\Delta$. Let $X_{v}$ be the random variable corresponding to the number of colours that are assigned to at least two vertices in $N_{v}$ and are retained by all these vertices. We also define $Y_{v}$ as the number of colours assigned to at least two vertices in $N_{v}$ that are not linked by a digon, but not necessarily retained. Finally, we define $Z_{v}$ as the number of colours assigned to at least two vertices in $N_{v}$ that are not linked by a digon and one of these vertices does not retain its colour. We thus have $X_{v}=Y_{v}-Z_{v}$.
We let $A_{v}$ be the event that $X_{v}<\ell$. Our goal is to show that none of the events $A_{v}$ occurs with non-zero probability. To do so, we will prove that, for each vertex $v$, the probability that $A_{v}$ occurs is small enough and then conclude by the Lovász Local Lemma (Lemma 2.1).
@@
For every vertex $v$, we have $\mathbf{E}(X_{v})\geq\frac{B_{v}}{2e^{7}\Delta}$
Proof of claim. For every pair of vertices $u,w\in N_{v}$ and every colour $c\in[k]$, we let $A_{u,w}^{c}$ be the event that $u,w$ are exactly the vertices coloured $c$ in $N_{v}$ and they both retain their colour. We let $X_{u,w}^{c}$ be the binary random variable equals to $1$ if $A_{u,w}^{c}$ occurs and 00 otherwise. We clearly have $X_{v}\geq\sum_{u,w,c}X_{u,w}^{c}$.
Assume $u$ and $w$ are not linked by a digon. If $uw$ is a simple arc of $D$, then $A_{u,w}^{c}$ occurs if and only if both $u$ and $w$ are coloured $c$ and no vertex in $N=(N_{v}\cup N^{-}(u)\cup N^{+}(w))\setminus\{u,w\}$ receives colour $c$. If the arc $uw$ is missing, then $A_{u,w}^{c}$ occurs in particular if both $u$ and $w$ are coloured $c$ and no vertex in $N=(N_{v}\cup N^{+}(u)\cup N^{-}(w))\setminus\{u,w\}$ receives colour $c$. In both cases, as $k=\left\lfloor\frac{\Delta}{2}\right\rfloor$, we have $|N|\leq 3\Delta\leq 7k$, which implies
$\mathbf{Pr}(A_{u,w}^{c})\geq\left(\frac{1}{k}\right)^{2}\left(1-\frac{1}{k}\right)^{7k}.$
@@
$\mathbf{E}(X_{v})\geq k\frac{B_{v}}{2}\left(\frac{1}{k}\right)^{2}\left(1-\frac{1}{k}\right)^{7k}\geq\frac{B_{v}}{\Delta}\left(1-\frac{1}{k}\right)^{7k}\geq\frac{B_{v}}{2e^{7}\Delta},$
where in the last inequality we used that $k=\left\lfloor\frac{\Delta}{2}\right\rfloor$ is large enough and ${\displaystyle\lim_{k\to+\infty}}\left(1-\frac{1}{k}\right)^{7k}=\frac{1}{e^{7}}>\frac{1}{2e^{7}}$$\lozenge$
@@
For every vertex $v$, we have $\mathbf{E}(X_{v})\leq\mathbf{E}(Y_{v})\leq\frac{3B_{v}}{\Delta}$
Proof of claim. Since $X_{v}=Y_{v}-Z_{v}$ it is clear that $\mathbf{E}(X_{v})\leq\mathbf{E}(Y_{v})$. Given two vertices $u,w\in N_{v}$ that are not linked by a digon and a colour $c\in[k]$, we let $Y_{u,w}^{c}$ be the binary random variable equals to $1$ if both $u$ and $w$ are coloured $c$ and 00 otherwise. We have $Y_{v}\leq\sum_{u,w,c}Y_{u,w}^{c}$ and $\mathbf{Pr}(Y_{u,w}^{c}=1)=\frac{1}{k^{2}}$. Since we have at most $B_{v}$ choices for the pair $u,w$ and exactly $k$ for the colour $c$, by linearity of the expectation we obtain
$\mathbf{E}(X_{v})\leq\mathbf{E}(Y_{v})\leq kB_{v}\left(\frac{1}{k}\right)^{2}\leq\frac{3B_{v}}{\Delta},$
where in the last inequality we used $k=\left\lfloor\frac{\Delta}{2}\right\rfloor\geq\frac{\Delta}{3}$. $\lozenge$
@@
For every vertex $v$, $\mathbf{Pr}\left(|X_{v}-\mathbf{E}(X_{v})|>\log\Delta\sqrt{\mathbf{E}(X_{v})}\right)<\frac{1}{\Delta^{5}}$.
Proof of claim. Let us set $t=\frac{1}{2}\log\Delta\sqrt{\mathbf{E}(X_{v})}$. Recall that $X_{v}=Y_{v}-Z_{v}$, so by linearity of the expectation we have $\mathbf{E}(X_{v})=\mathbf{E}(Y_{v})-\mathbf{E}(Z_{v})$, and $|X_{v}-\mathbf{E}(X_{v})|\leq|Y_{v}-\mathbf{E}(Y_{v})|+|Z_{v}-\mathbf{E}(Z_{v})|$. In particular, this implies
$\mathbf{Pr}(|X_{v}-\mathbf{E}(X_{v})|>2t)\leq\mathbf{Pr}(|Y_{v}-\mathbf{E}(Y_{v})|>t)+\mathbf{Pr}(|Z_{v}-\mathbf{E}(Z_{v})|>t).$ (1)
We will use Talagrand’s inequality (Lemma 2.2) to bound both $\mathbf{Pr}(|Y_{v}-\mathbf{E}(Y_{v})|>t)$ and $\mathbf{Pr}(|Z_{v}-\mathbf{E}(Z_{v})|>t)$. Observe that changing the colour of exactly one vertex affects $Y_{v}$ by at most $1$. Also if $Y_{v}\geq s$ then there exists a set of $2s$ vertices whose colours certify $Y_{v}\geq s$: for each of the $s$ colours we may choose two vertices not linked by a digon sharing this colour. Combining Claims 3.5 and 3.6 we obtain $\mathbf{E}(Y_{v})\leq 6e^{7}\mathbf{E}(X_{v})$. Hence, $\Delta$ being is sufficiently large, we have $t>96\sqrt{2\mathbf{E}(Y_{v})}+256$, which allows us to apply Talagrand’s inequality to $Y_{v}$ and obtain
$\mathbf{Pr}(|Y_{v}-\mathbf{E}(Y_{v})|>t)\leq 4\exp\left(-\frac{t^{2}}{64\mathbf{E}(Y_{v})+16t}\right).$ (2)
We now consider $Z_{v}$. Changing the colour of exactly one vertex affects $Z_{v}$ by at most $2$. Also if $Z_{v}\geq s$ then there exists a set of $4s$ vertices whose colours certify $Z_{v}\geq s$: for each of the $s$ colours we may choose four vertices $a,b,b^{-},b^{+}$ sharing this colour, $a\neq b$, such that $a$ and $b$ are not linked by a digon in $D$, and $b^{-}b,bb^{+}\in A(D)$ ($a,b^{-},b^{+}$ are not necessarily distinct). Observe that $Z_{v}\leq Y_{v}$, so in particular $\mathbf{E}(Z_{v})\leq\mathbf{E}(Y_{v})\leq 6e^{7}\mathbf{E}(X_{v})$ by Claims 3.5 and 3.6. Hence, $\Delta$ being sufficiently large, we have $t>192\sqrt{4\mathbf{E}(Z_{v})}+2048$, which allows us to apply Talagrand’s inequality to $Z_{v}$ and obtain
$\mathbf{Pr}(|Z_{v}-\mathbf{E}(Z_{v})|>t)\leq 4\exp\left(-\frac{t^{2}}{512\mathbf{E}(Z_{v})+128t}\right).$ (3)
Combining (1), (2), and (3) together with $\mathbf{E}(Z_{v})\leq\mathbf{E}(Y_{v})\leq 6e^{7}\mathbf{E}(X_{v})$ we obtain
$\mathbf{Pr}(|X_{v}-\mathbf{E}(X_{v})|>2t)$ $\leq 8\exp\left(-\frac{t^{2}}{3072e^{7}\mathbf{E}(X_{v})+128t}\right)=8\exp\left(-\frac{\log^{2}\Delta}{12288e^{7}+256\frac{\log\Delta}{\sqrt{\mathbf{E}(X_{v})}}}\right)$
$\leq 8\exp\left(-\frac{\log^{2}\Delta}{12288e^{7}+256\sqrt{2e^{7}}\frac{1}{\sqrt{\log\Delta}}}\right),$
where in the last inequality we used $\mathbf{E}(X)\geq\frac{B_{v}}{2e^{7}\Delta}$ and $B_{v}\geq\Delta\log^{3}\Delta$. As $\Delta$ is sufficiently large, we obtain
$\mathbf{Pr}\left(|X_{v}-\mathbf{E}(X_{v})|>\log\Delta\sqrt{\mathbf{E}(X_{v})}\right)<\frac{1}{\Delta^{5}},$
@@
By Claim 3.5 together with the fact that $B_{v}\geq B\geq\Delta\log^{3}\Delta$, we have
$\mathbf{E}(X_{v})-\log\Delta\sqrt{\mathbf{E}(X_{v})}\geq\frac{B_{v}}{2e^{7}\Delta}\left(1-\log\Delta\sqrt{\frac{2e^{7}\Delta}{B_{v}}}\right)\geq\frac{B}{2e^{7}\Delta}\left(1-\sqrt{\frac{2e^{7}}{\log\Delta}}\right).$
By choice of $\ell$, we thus have $\mathbf{E}(X_{v})-\log\Delta\sqrt{\mathbf{E}(X_{v})}\geq 2\ell\left(1-\sqrt{\frac{2e^{7}}{\log\Delta}}\right)$, which is larger than $\ell$ for $\Delta$ large enough. Hence we obtain
$\mathbf{Pr}(A_{v})=\mathbf{Pr}\left(X_{v}<\ell\right)\leq\mathbf{Pr}\left(X_{v}<\mathbf{E}(X)-\log\Delta\sqrt{\mathbf{E}(X)}\right)\leq\mathbf{Pr}\left(|X_{v}-\mathbf{E}(X)|>\log\Delta\sqrt{\mathbf{E}(X)}\right).$
Therefore, by Claim 3.7, we have $\mathbf{Pr}(A_{v})<\frac{1}{\Delta^{5}}$. Observe finally that the value of $X_{v}$ depends only on the colour of vertices at distance at most $2$ of $v$ in $\operatorname{UG}(D)$. Hence, each event $A_{v}$ is mutually independent of all other events $A_{u}$ except at most $(2\Delta)^{4}=d$ of them. For every vertex $v$, we have $e\mathbf{Pr}(A_{v})(d+1)\leq e\frac{16\Delta^{4}+1}{\Delta^{5}}$, which is smaller than $1$ if $\Delta$ is large enough. Hence by the Lovász Local Lemma (Lemma 2.1) none of the events $A_{v}$ occurs with non-zero probability, which concludes the proof. ∎

### v3

@@ 4 Colouring sparse digraphs
In view of proving Theorem 1.7, our first goal is to generalise the following result due to Molloy and Reed [30, Theorem 10.5] on the chromatic number of sparse graphs. For a vertex $v$ of a graph $G$, we denote by $m(v)$ the number of edges of $G[N(v)]$. We say that $G$ is $B$-sparse if $m(v)\leqslant\binom{\Delta(G)}{2}-B$ for every $v\in V(G)$.
@@
$\chi(G)\leqslant(\Delta+1)-\frac{B}{e^{6}\Delta}.$
@@
There exists $\Delta_{1}\geqslant 2$ such that the following holds. Let $D$ be a digraph with maximum degree $\Delta_{\max}(D)=\Delta\geqslant\Delta_{1}$ and $a\in\mathbb{R}$ be such that $a>\frac{\log^{3}\Delta}{\Delta-1}$, and for every vertex $v\in V(D)$, $\min(m^{+}(v),m^{-}(v))\leqslant(1-a)\Delta(\Delta-1)$. Then
$\vec{\chi}(D)\leqslant\left(1-\frac{a}{5e^{7}}\right)\left(\Delta+1\right).$
@@
By assumption $D$ is $B_{a}$-sparse for $B_{a}=a\Delta(\Delta-1)>\Delta\log^{3}\Delta$. Hence, by taking $\Delta_{1}\geqslant\Delta_{0}$, we obtain by Theorem 1.8 that
$\vec{\chi}(D)\leqslant\left(1-\frac{a}{4e^{7}}\right)\left(\Delta+1\right)+\frac{a}{2e^{7}}\leqslant\left(1-\frac{a}{5e^{7}}\right)\left(\Delta+1\right),$
where in the second inequality we used $\Delta_{1}\geqslant 9$. ∎
@@
Let $D$ be a $B$-sparse digraph with maximum degree $\Delta_{\max}(D)=\Delta$. There exists a $B$-sparse $\Delta$-diregular digraph $H$ such that $\vec{\chi}(D)\leqslant\vec{\chi}(H)$.
@@
We repeat the following process until $D$ is $\Delta$-diregular. Add to $D$ a disjoint copy of $\vec{\reflectbox{$D$}}$ , the digraph obtained from $D$ by reversing every arc. Then for every vertex $v\in V(D)$ such that $d^{+}_{D}(v)<\Delta$, we add the arc $vv^{\prime}$, where $v^{\prime}$ is the copy of $v$ in $\vec{\reflectbox{$D$}}$ . Similarly, for every vertex $v$ such that $d^{-}_{D}(v)<\Delta$, we add the arc $v^{\prime}v$. Let $D^{\prime}$ be the obtained digraph. Note that $\vec{\chi}(D)\leqslant\vec{\chi}(D^{\prime})$ trivially holds as $D$ is a subdigraph of $D^{\prime}$. Observe also that, for every vertex $v\in V(D)$, we have $m^{+}_{D}(v)=m^{+}_{D^{\prime}}(v)=m^{-}_{D^{\prime}}(v^{\prime})$ and $m^{-}_{D}(v)=m^{-}_{D^{\prime}}(v)=m^{+}_{D^{\prime}}(v^{\prime})$. Hence, as $D$ is $B$-sparse, so is $D^{\prime}$.
@@
$\left|\{c\in[k]\mid\exists u,w\in N^{-}(v),u\neq w,\alpha(u)=\alpha(w)=c\}\right|\geqslant\ell$
or $\left|\{c\in[k]\mid\exists u,w\in N^{+}(v),u\neq w,\alpha(u)=\alpha(w)=c\}\right|\geqslant\ell.$
@@
Let $D$ be a $\Delta$-diregular digraph and $k,\ell$ be two integers such that $k\leqslant\Delta+1$-$\ell$. If $D$ admits a partial $(k,\ell)$-dicolouring, then $\vec{\chi}(D)\leqslant\Delta+1$-$\ell$.
@@
Assume that $D$ admits a partial $(k,\ell)$-dicolouring. We let $\alpha$ be a partial $(\Delta+1-\ell,\ell)$-dicolouring of $D$ for which the number of uncoloured vertices is minimum, the existence of which is guaranteed as $k\leqslant\Delta+1$-$\ell$. We claim that $\alpha$ is a dicolouring of $D$, hence implying $\vec{\chi}(D)\leqslant\Delta+1$-$\ell$. Assume this is not the case, and there is a vertex $v$ which is not coloured by $\alpha$. By definition of a partial $(\Delta+1-\ell,\ell)$-dicolouring, we have $|\alpha(N^{-}(v))|\leqslant\Delta$-$\ell$ or $|\alpha(N^{+}(v))|\leqslant\Delta$-$\ell$. Hence we can choose for $v$ a colour of $\Delta+1-\ell]$ that is not appearing either in its in- or in its out-neighbourhood. Colouring $v$ with this specific colour extends $\alpha$ into a partial $(\Delta+1-\ell,\ell)$-dicolouring of $D$ with fewer uncoloured vertices, a contradiction to the choice of $\alpha$. ∎
@@
We do not give the explicit value of $\Delta_{0}$, we simply assume in what follows that it is arbitrarily large. We assume that $D$ is $\Delta$-diregular, for otherwise we may prove the result for the $B$-sparse $\Delta$-diregular digraph $H$ obtained by Lemma 4.3, and conclude that $\vec{\chi}(D)\leqslant\vec{\chi}(H)\leqslant(\Delta+1)-\frac{B}{4e^{7}\Delta}$.
We use the probabilistic method. Let us fix $k=\left\lfloor\frac{\Delta}{2}\right\rfloor$. We randomly colour every vertex of $D$ with a colour of $[k]$. Then we simultaneously uncolour every vertex $v$ having at least one in-neighbour and at least one out-neighbour sharing its colour. Our goal is to show that with non-zero probability, the obtained colouring $\alpha$ is a partial $(k,\ell)$-dicolouring of $D$, where $\ell=\frac{B}{4e^{7}\Delta}$. Observe that we have $k\leqslant\Delta+1$-$\ell$ as $B\leqslant\Delta(\Delta-1)$. Hence, if $\alpha$ is indeed a partial $(k,\ell)$-dicolouring of $D$, the result follows by Lemma 4.4. Note that $\alpha$ is necessarily a dicolouring of an induced subdigraph of $D$, for otherwise $D$, partially coloured with $\alpha$, contains a monochromatic directed cycle, the vertices of which should have been uncoloured in the process. It remains to show that, for every vertex of $D$, $\ell$ colours are repeated either in its in-neighbourhood or in its out-neighbourhood.
For each vertex $v$, we let $N_{v}$ be $N^{+}(v)$ if $m^{+}(v)\leqslant m^{-}(v)$ and $N^{-}(v)$ otherwise. We also let $B_{v}$ be the number of missing arcs in $D[N_{v}]$, that is $B_{v}=\Delta(\Delta-1)-\min(m^{-}(v),m^{+}(v))$. Note that $B_{v}\geqslant B>\Delta\log^{3}\Delta$. Let $X_{v}$ be the random variable corresponding to the number of colours that are assigned to at least two vertices in $N_{v}$ and are retained by all these vertices. We also define $Y_{v}$ as the number of colours assigned to at least two vertices in $N_{v}$ that are not linked by a digon, but not necessarily retained. Finally, we define $Z_{v}$ as the number of colours assigned to at least two vertices in $N_{v}$ that are not linked by a digon and one of these vertices does not retain its colour. We thus have $X_{v}=Y_{v}-Z_{v}$.
We let $A_{v}$ be the event that $X_{v}<\ell$. Our goal is to show that none of the events $A_{v}$ occur with non-zero probability. To do so, we will prove that, for each vertex $v$, the probability that $A_{v}$ occurs is small enough and then conclude by the Lovász Local Lemma (Lemma 2.1).
@@
For every vertex $v$, we have $\mathbb{E}(X_{v})\geqslant\frac{B_{v}}{2e^{7}\Delta}$
Proof of claim. For every pair of vertices $u,w\in N_{v}$ and every colour $c\in[k]$, we let $A_{u,w}^{c}$ be the event that $u,w$ are exactly the vertices coloured $c$ in $N_{v}$ and they both retain their colour. We let $X_{u,w}^{c}$ be the binary random variable equals to $1$ if $A_{u,w}^{c}$ occurs and 0 otherwise. We clearly have $X_{v}\geqslant\sum_{u,w,c}X_{u,w}^{c}$.
Assume $u$ and $w$ are not linked by a digon. If $uw$ is a simple arc of $D$, then $A_{u,w}^{c}$ occurs if and only if both $u$ and $w$ are coloured $c$ and no vertex in $N=(N_{v}\cup N^{-}(u)\cup N^{+}(w))\setminus\{u,w\}$ receives colour $c$. If the arc $uw$ is missing, then $A_{u,w}^{c}$ occurs in particular if both $u$ and $w$ are coloured $c$ and no vertex in $N=(N_{v}\cup N^{+}(u)\cup N^{-}(w))\setminus\{u,w\}$ receives colour $c$. In both cases, as $k=\left\lfloor\frac{\Delta}{2}\right\rfloor$, we have $|N|\leqslant 3\Delta\leqslant 7k$, which implies
$\mathbb{P}(A_{u,w}^{c})\geqslant\left(\frac{1}{k}\right)^{2}\left(1-\frac{1}{k}\right)^{7k}.$
@@
$\mathbb{E}(X_{v})\geqslant k\frac{B_{v}}{2}\left(\frac{1}{k}\right)^{2}\left(1-\frac{1}{k}\right)^{7k}\geqslant\frac{B_{v}}{\Delta}\left(1-\frac{1}{k}\right)^{7k}\geqslant\frac{B_{v}}{2e^{7}\Delta},$
where in the last inequality we used that $k=\left\lfloor\frac{\Delta}{2}\right\rfloor$ is large enough and that ${\displaystyle\lim_{k\to+\infty}}\left(1-\frac{1}{k}\right)^{7k}=\frac{1}{e^{7}}>\frac{1}{2e^{7}}\lozenge$
@@
For every vertex $v$, we have $\mathbb{E}(X_{v})\leqslant\mathbb{E}(Y_{v})\leqslant\frac{3B_{v}}{\Delta}$
Proof of claim. Since $X_{v}=Y_{v}-Z_{v}$ it is clear that $\mathbb{E}(X_{v})\leqslant\mathbb{E}(Y_{v})$. Given two vertices $u,w\in N_{v}$ that are not linked by a digon and a colour $c\in[k]$, we let $Y_{u,w}^{c}$ be the binary random variable equals to $1$ if both $u$ and $w$ are coloured $c$ and 0 otherwise. We have $Y_{v}\leqslant\sum_{u,w,c}Y_{u,w}^{c}$ and $\mathbb{P}(Y_{u,w}^{c}=1)=\frac{1}{k^{2}}$. Since we have at most $B_{v}$ choices for the pair $u,w$ and exactly $k$ for the colour $c$, by linearity of the expectation we obtain
$\mathbb{E}(X_{v})\leqslant\mathbb{E}(Y_{v})\leqslant kB_{v}\left(\frac{1}{k}\right)^{2}\leqslant\frac{3B_{v}}{\Delta},$
where in the last inequality we used $k=\left\lfloor\frac{\Delta}{2}\right\rfloor\geqslant\frac{\Delta}{3}$. $\lozenge$
@@
For every vertex $v$, $\mathbb{P}\left(|X_{v}-\mathbb{E}(X_{v})|>\log\Delta\sqrt{\mathbb{E}(X_{v})}\right)<\frac{1}{\Delta^{5}}$.
Proof of claim. Let us set $t=\frac{1}{2}\log\Delta\sqrt{\mathbb{E}(X_{v})}$. Recall that $X_{v}=Y_{v}-Z_{v}$, so by linearity of the expectation we have $\mathbb{E}(X_{v})=\mathbb{E}(Y_{v})-\mathbb{E}(Z_{v})$, and $|X_{v}-\mathbb{E}(X_{v})|\leqslant|Y_{v}-\mathbb{E}(Y_{v})|+|Z_{v}-\mathbb{E}(Z_{v})|$. In particular, this implies
$\mathbb{P}(|X_{v}-\mathbb{E}(X_{v})|>2t)\leqslant\mathbb{P}(|Y_{v}-\mathbb{E}(Y_{v})|>t)+\mathbb{P}(|Z_{v}-\mathbb{E}(Z_{v})|>t).$ (2)
We will use Talagrand’s inequality (Lemma 2.2) to bound both $\mathbb{P}(|Y_{v}-\mathbb{E}(Y_{v})|>t)$ and $\mathbb{P}(|Z_{v}-\mathbb{E}(Z_{v})|>t)$. Observe that changing the colour of exactly one vertex affects $Y_{v}$ by at most $1$. Also if $Y_{v}\geqslant s$ then there exists a set of $2s$ vertices whose colours certify $Y_{v}\geqslant s$: for each of the $s$ colours we may choose two vertices not linked by a digon sharing this colour. Combining Claims 4.5 and 4.6 we obtain $\mathbb{E}(Y_{v})\leqslant 6e^{7}\mathbb{E}(X_{v})$. Hence, $\Delta$ being is sufficiently large, we have $t>126\sqrt{2\mathbb{E}(Y_{v})}+688$, which allows us to apply Talagrand’s inequality to $Y_{v}$ and obtain
$\mathbb{P}(|Y_{v}-\mathbb{E}(Y_{v})|>t)\leqslant 4\exp\left(-\frac{t^{2}}{64(\mathbb{E}(Y_{v})+t)}\right).$ (3)
We now consider $Z_{v}$. Changing the colour of exactly one vertex affects $Z_{v}$ by at most $1$. Also if $Z_{v}\geqslant s$ then there exists a set of $4s$ vertices whose colours certify $Z_{v}\geqslant s$: for each of the $s$ colours we may choose four vertices $a,b,b^{-},b^{+}$ sharing this colour, $a\neq b$, such that $a$ and $b$ are not linked by a digon in $D$, and $b^{-}b,bb^{+}\in A(D)$ ($a,b^{-},b^{+}$ are not necessarily distinct). Observe that $Z_{v}\leqslant Y_{v}$, so in particular $\mathbb{E}(Z_{v})\leqslant\mathbb{E}(Y_{v})\leqslant 6e^{7}\mathbb{E}(X_{v})$ by Claims 4.5 and 4.6. Hence, $\Delta$ being sufficiently large, we have $t>126\sqrt{4\mathbb{E}(Z_{v})}+1376$, which allows us to apply Talagrand’s inequality to $Z_{v}$ and obtain
$\mathbb{P}(|Z_{v}-\mathbb{E}(Z_{v})|>t)\leqslant 4\exp\left(-\frac{t^{2}}{128(\mathbb{E}(Z_{v})+t)}\right).$ (4)
Combining (2), (3), and (4) together with $\mathbb{E}(Z_{v})\leqslant\mathbb{E}(Y_{v})\leqslant 6e^{7}\mathbb{E}(X_{v})$ we obtain
$\mathbb{P}(|X_{v}-\mathbb{E}(X_{v})|>2t)$ $\leqslant 8\exp\left(-\frac{t^{2}}{768e^{7}\mathbb{E}(X_{v})+128t}\right)=8\exp\left(-\frac{\log^{2}\Delta}{3072e^{7}+256\frac{\log\Delta}{\sqrt{\mathbb{E}(X_{v})}}}\right)$
$\leqslant 8\exp\left(-\frac{\log^{2}\Delta}{3072e^{7}+256\sqrt{2e^{7}}\frac{1}{\sqrt{\log\Delta}}}\right),$
where in the last inequality we used $\mathbb{E}(X_{v})\geqslant\frac{B_{v}}{2e^{7}\Delta}$ and $B_{v}\geqslant\Delta\log^{3}\Delta$. As $\Delta$ is sufficiently large, we obtain
$\mathbb{P}\left(|X_{v}-\mathbb{E}(X_{v})|>\log\Delta\sqrt{\mathbb{E}(X_{v})}\right)<\frac{1}{\Delta^{5}},$
@@
By Claim 4.5 together with the fact that $B_{v}\geqslant B\geqslant\Delta\log^{3}\Delta$, we have
$\mathbb{E}(X_{v})-\log\Delta\sqrt{\mathbb{E}(X_{v})}\geqslant\frac{B_{v}}{2e^{7}\Delta}\left(1-\log\Delta\sqrt{\frac{2e^{7}\Delta}{B_{v}}}\right)\geqslant\frac{B}{2e^{7}\Delta}\left(1-\sqrt{\frac{2e^{7}}{\log\Delta}}\right).$
By choice of $\ell$, we thus have $\mathbb{E}(X_{v})-\log\Delta\sqrt{\mathbb{E}(X_{v})}\geqslant 2\ell\left(1-\sqrt{\frac{2e^{7}}{\log\Delta}}\right)$, which is larger than $\ell$ for $\Delta$ large enough. Hence we obtain
$\mathbb{P}(A_{v})=\mathbb{P}\left(X_{v}<\ell\right)\leqslant\mathbb{P}\left(X_{v}<\mathbb{E}(X)-\log\Delta\sqrt{\mathbb{E}(X)}\right)\leqslant\mathbb{P}\left(|X_{v}-\mathbb{E}(X)|>\log\Delta\sqrt{\mathbb{E}(X)}\right).$
Therefore, by Claim 4.7, we have $\mathbb{P}(A_{v})<\frac{1}{\Delta^{5}}$. Observe finally that the value of $X_{v}$ depends only on the colour of vertices at distance at most $2$ of $v$ in $\operatorname{UG}(D)$. Hence, each event $A_{v}$ is mutually independent of all other events $A_{u}$ except at most $(2\Delta)^{4}=d$ of them. For every vertex $v$, we have $e\mathbb{P}(A_{v})(d+1)\leqslant e\frac{16\Delta^{4}+1}{\Delta^{5}}$, which is smaller than $1$ if $\Delta$ is large enough. Hence by the Lovász Local Lemma (Lemma 2.1) none of the events $A_{v}$ occurs with non-zero probability, which concludes the proof. ∎